# Online LLM watermark detection via e-processes

### v1

@@ Abstract
Watermarking for large language models (LLMs) has emerged as an effective tool for distinguishing AI-generated text from human-written content. Statistically, watermark schemes induce dependence between generated tokens and a pseudo-random sequence, reducing watermark detection to a hypothesis testing problem on independence. We develop a unified framework for LLM watermark detection based on e-processes, providing anytime-valid guarantees for online testing. We propose various methods to construct empirically adaptive e-processes that can enhance the detection power. In addition, theoretical results are established to characterize the power properties of the proposed procedures. Some experiments demonstrate that the proposed framework achieves competitive performance compared to existing watermark detection methods.
@@ 1 Introduction
The recent advent of powerful Large Language Models (LLMs) has fundamentally transformed natural language generation. State-of-the-art LLMs such as GPT-4 (Achiam et al.,, 2024) and LLaMA (Grattafiori et al.,, 2024) demonstrate remarkable capabilities in producing human-like text. While these advances unlock substantial societal and economic benefits, they also raise significant ethical and security concerns. The increasing difficulty in distinguishing AI-generated text from human-written content facilitates its misuse for malicious purposes, including generation of misinformation (Zellers et al.,, 2019; Starbird,, 2019; Weidinger et al.,, 2021), academic plagiarism (Stokel-Walker,, 2022; Milano et al.,, 2023), and automated fraudulent content generation (Shumailov et al.,, 2023; Das et al.,, 2024). Such misuse may seriously erode trust in digital ecosystems. To mitigate these problems, there is a pressing need for effective mechanisms to make AI-generated text distinguishable while preserving the quality and usability of language models.
@@
Watermarking has emerged as a promising approach to address this challenge (Fernandez et al.,, 2023; Kuditipudi et al.,, 2024; Hu et al.,, 2024). In watermark schemes for LLMs, a subtle algorithmic signal is embedded into the text generation process, enabling post-hoc verification of AI authorship without significantly degrading text quality. From a statistical perspective, watermark schemes modify the sampling procedure of the language model.
@@
Concretely, let $\mathcal{W}$ denote the vocabulary of tokens. Given a sequence of previous tokens $(w_{1},\ldots,w_{t-1})$, an unwatermarked language model computes a next-token prediction (NTP) distribution $\mathbf{P}_{t}$ and samples $w_{t}$ from $\mathbf{P}_{t}$. In contrast, a watermarked model generates $w_{t}$ according to both $\mathbf{P}_{t}$ and a pseudo-random variable (or vector) $\zeta_{t}$. In practice, the pseudo-random sequence can be reconstructed using the generated text together with shared secret information, referred to as the watermark key. A formal description of this mechanism is deferred to Section 3. Crucially, the watermark mechanism induces statistical dependence between $w_{t}$ and $\zeta_{t}$. In unwatermarked text generation, $w_{t}$ is independent of $\zeta_{t}$, whereas in watermarked generation, this independence is deliberately violated. Therefore, watermark detection can be formulated as a hypothesis testing problem on the dependence structure between the observed tokens and the pseudo-random variables (Aaronson,, 2023; Li et al., 2025b, ; Li et al., 2025c, ; Piet et al.,, 2025).
Despite substantial progress, existing watermark detection methods still face several challenges. First, most procedures are designed for fixed-sample testing and lack rigorous guarantees under optional stopping, which is particularly problematic in sequential or streaming settings. Second, statistical power may deteriorate when the NTP distribution becomes highly concentrated (i.e., nearly degenerate), a phenomenon that may occur in long text generation. Third, theoretical characterizations of testing power remain limited in general watermark detection frameworks.
@@
To address these issues, we develop a unified framework for watermark detection based on e-values and e-processes, which are recently developed tools for online testing with useful features (Grünwald et al.,, 2024; Vovk and Wang,, 2021; Wang and Ramdas,, 2022; Ramdas et al.,, 2023); see Ramdas and Wang, (2025) for background and a comprehensive treatment of the topic. Our framework reformulates watermark detection as sequential hypothesis testing of independence and constructs e-processes tailored to the watermark-induced structure, with a special focus on the popular Gumbel-max watermark (Aaronson,, 2023). The proposed approach enjoys the following key advantages:
@@
    By using adaptive weights and the online Grenander algorithm (Hore et al.,, 2026) to construct calibrators, the proposed e-processes are flexible and have good performance. In some situations, they can even have better performance than the existing methods that do not have sequential validity.
@@
Together, these contributions provide a principled and theoretically grounded approach to watermark detection. Our framework unifies existing ideas under the lens of e-processes, clarifies the statistical structure of watermark detection, and offers both finite-sample guarantees and asymptotic efficiency. Empirical studies further demonstrate the practical advantages of the proposed methods.
@@ 3.1 Watermarks and the detection problem
and $\mathbb{Q}$ is an objective data generating probability (Li et al., 2025b, ). The condition in (1) can be rewritten as $S(\mathbf{P},\zeta)\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{d}}}\mathrm{Cat}(\mathbf{P})$ for each $\mathbf{P}\in\Delta_{K}$, where $\mathrm{Cat}(\mathbf{P})$ is the categorical distribution with probability vector $\mathbf{P}$. There is no other requirement on $S$.
@@
For a given $\mathbf{P}$, we use $\mathbb{U}^{\mathbf{P}}$ for the distribution of $(W,\zeta)$ under $H_{0}$ and $\mathbb{Q}^{\mathbf{P}}$ for the distribution of $(W,\zeta)$ under $H_{1}$, to emphasize that both $\mathbb{U}^{\mathbf{P}}$ and $\mathbb{Q}^{\mathbf{P}}$ depends on the usually unobservable $\mathbf{P}$. This makes the testing problem consisting of composite null and composite alternative. Moreover, $\mathbb{Q}^{\mathbf{P}}$ depends on the watermarking scheme $S$ through the condition (1) on $\mathbb{Q}^{\mathbf{P}}$.
@@ 3.3 Static information bounds
and it is $\infty$ otherwise. By Ramdas and Wang, (2025, Theorem 3.20),
@@
it is necessary and sufficient that $E$ is given by $E=\mathrm{d}\mathbb{Q}^{\mathbf{P}}/\mathrm{d}\mathbb{U}^{\mathbf{P}}$ (Ramdas and Wang,, 2025, Theorem 3.20), which depends on the NTP vector $\mathbf{P}$. Moreover, this is usually not a feasible choice of an e-variable, because the constraint in (2),
@@ 3.4 The Gumbel-max watermark
We next describe the Gumbel-max watermark, which is proposed by Aaronson, (2023). As one of the most influential watermark schemes, it has been implemented internally at OpenAI (OpenAI,, 2024) and has become a baseline in many studies (Li et al., 2025b, ; Li et al., 2025c, ). This watermark utilizes the Gumbel-max trick (Gumbel,, 1954; Jang et al.,, 2017) to ensure the unbiasedness characterized in (1). The Gumbel-max watermark operates with the pseudo-random vector $\zeta=(U_{w})_{w\in\mathcal{W}}$ consisting of iid standard uniform variables, and the decoder is given by
@@
As suggested by Aaronson, (2023), the pivotal statistic for one token is
@@
Putting the subscript $t$ back to the original detection problem for $T<\infty$ tokens, we denote by $Y_{t}=U_{t,W_{t}}$ the pivotal statistic for the $t$-th token $W_{t}$. Aaronson, (2023) proposed to declare the presence of the watermark if the sum-based statistic $H_{\mathrm{ars}}=\sum_{t=1}^{T}h_{\mathrm{ars}}(Y_{t})$, with the score function $h_{\mathrm{ars}}(y)=-\log(1-y)$, is above a certain threshold. Other score functions, such as the log function $h_{\mathrm{log}}(y)=\log y$ (Fernandez et al.,, 2023; Kuditipudi et al.,, 2024), and the optimized score function $h_{\mathrm{gum}}^{*}$ analyzed in Li et al., 2025b , can also be adopted in the sum-based approach.
@@
for $x,y\in A$. We refer to Marshall et al., (2011) for a treatment of majorization and Schur-convexity.
@@
By Proposition C.1 of Marshall et al., (2011), the sum of univariate convex functions is Schur-convex. Therefore, for statement (i), it suffices to prove that $p\mapsto py^{1/p}$ is convex on $[0,1]$, which follows by straightforward calculus. For statement (ii), it suffices to note that $\mathbf{P}\preccurlyeq\mathbf{P}^{*}$ for all $\mathbf{P}\in\Delta^{\delta}_{K}$, which can be checked by definition. Statement (iii) follows directly from $F^{\mathbf{P}}(y)\leq F^{\mathbf{P}^{*}}(y)$ for any $y\in[0,1]$ and $\mathbf{P}\in\Delta_{K}^{\delta}$, as shown in (ii). ∎
Using the Gumbel-max watermark, the maximum e-power for an e-variable is the KL divergence between the two distributions (Theorem 3.20 of Ramdas and Wang,, 2025), that is,
@@ 4 E-processes for watermarks
Throughout this section, we let $Y_{t},~t\in\mathbb{T}$ be the pivotal statistics for a watermark scheme. Recall that $\mathcal{F}=(F_{t})_{t\in\mathbb{T}_{0}}$ is the information filtration, to which $(Y_{t})_{t\in\mathbb{T}}$ is adapted, with the interpretation that $\mathcal{F}_{t-1}$ contains all the statistical information available before generating the token $w_{t}$. The sequence $(Y_{t})_{t\in\mathbb{T}}$ satisfies two basic conditions.
@@
    $Y_{t}$ is uniformly distributed on $[0,1]$ and independent of $\mathcal{F}_{t-1}$ under the null hypothesis.
@@
    $Y_{t}$ is super-uniform conditional on $\mathcal{F}_{t-1}$ under the alternative hypothesis.
@@ 4.1 Online test with e-processes
We first introduce (p-to-e) calibrators (Vovk and Wang,, 2021), which convert p-values into e-values. A calibrator is a decreasing function $f:[0,1]\to\mathbb{R}_{+}$ satisfying $\int_{0}^{1}f(p)\mathrm{d}p=1$. Although sometimes only $\int_{0}^{1}f(p)\mathrm{d}p\leq 1$ is required, it is without loss of generality to consider only those with $\int_{0}^{1}f(p)\mathrm{d}p=1$.
@@
where each $E_{t}=f_{t}(1-Y_{t})$ for some $f_{t}\in\Phi_{t}$. We state the following basic result, which follows from standard arguments in the e-value literature. We present it because it is the most important technical tool for constructing our tests.
@@
Under Assumption 1 (i), the process $M=(M_{t})_{t\in\mathbb{T}_{0}}$ in (9) with
@@
is an e-process (also a supermartingale), where $f_{t}\in\Phi_{t}$ for each $t$. In particular, under the null hypothesis,
@@
$\mathbb{E}^{\mathbb{P}}[E_{t}|\mathcal{F}_{t-1}]=\mathbb{E}^{\mathbb{P}}[f_{t}(1-Y_{t})|\mathcal{F}_{t-1}]=\int_{0}^{1}f_{t}(p)\mathrm{d}p\leq 1,$
@@
where we have used that $Y_{t}$ is independent of $\mathcal{F}_{t-1}$ and that $f_{t}$ is determined by $Y_{1},\dots,Y_{t-1}$ and it integrates to at most $1$. Moreover, it is clear that $E_{t}$ is nonnegative. Therefore,
$\mathbb{E}^{\mathbb{P}}[M_{t}|\mathcal{F}_{t-1}]\leq M_{t-1},~~~t\in\mathbb{T}.$
@@
Thus, $M$ is a nonnegative supermartingale starting at $1$ and hence an e-process. The last standard statement follows from Ville’s inequality or the optional stopping theorem. ∎
@@
In the next two sections, we present two different ways of choosing the calibrator $f_{t}\in\Phi_{t}$ adaptively.

### v2

@@ Abstract
Watermarking for large language models (LLMs) has emerged as an effective tool for distinguishing AI-generated text from human-written content. Statistically, watermark schemes induce dependence between generated tokens and a pseudo-random sequence, reducing watermark detection to a hypothesis testing problem on independence. We develop a unified framework for LLM watermark detection based on e-processes, providing anytime-valid guarantees for online testing. We propose various methods to construct empirically adaptive e-processes that can enhance the detection power. The proposed methods are applicable to any sequential testing problem where independent pivotal statistics are available. In addition, theoretical results are established to characterize the power properties of the proposed procedures. Some experiments demonstrate that the proposed framework achieves competitive performance compared to existing watermark detection methods.
@@ 1 Introduction
The recent advent of powerful Large Language Models (LLMs) has fundamentally transformed natural language generation. State-of-the-art LLMs such as GPT-4 (Achiam et al., 2024) and LLaMA (Grattafiori et al., 2024) demonstrate remarkable capabilities in producing human-like text. While these advances unlock substantial societal and economic benefits, they also raise significant ethical and security concerns. The increasing difficulty in distinguishing AI-generated text from human-written content facilitates its misuse for malicious purposes, including generation of misinformation (Zellers et al., 2019; Starbird, 2019; Weidinger et al., 2021), academic plagiarism (Stokel-Walker, 2022; Milano et al., 2023), and automated fraudulent content generation (Shumailov et al., 2023; Das et al., 2024). Such misuse may seriously erode trust in digital ecosystems. To mitigate these problems, there is a pressing need for effective mechanisms to make AI-generated text distinguishable while preserving the quality and usability of language models.
@@
Watermarking has emerged as a promising approach to address this challenge (Fernandez et al., 2023; Kuditipudi et al., 2024; Hu et al., 2024). In watermark schemes for LLMs, a subtle algorithmic signal is embedded into the text generation process, enabling post-hoc verification of AI authorship without significantly degrading text quality. From a statistical perspective, watermark schemes modify the sampling procedure of the language model.
@@
Concretely, let $\mathcal{W}$ denote the vocabulary of tokens. Given a sequence of previous tokens $(w_{1},\ldots,w_{t-1})$, an unwatermarked language model computes a next-token prediction (NTP) distribution $\mathbf{P}_{t}$ and samples $w_{t}$ from $\mathbf{P}_{t}$. In contrast, a watermarked model generates $w_{t}$ according to both $\mathbf{P}_{t}$ and a pseudo-random variable (or vector) $\zeta_{t}$. In practice, the pseudo-random sequence can be reconstructed using the generated text together with shared secret information, referred to as the watermark key. A formal description of this mechanism is deferred to Section 3. Crucially, the watermark mechanism induces statistical dependence between $w_{t}$ and $\zeta_{t}$. In unwatermarked text generation, $w_{t}$ is independent of $\zeta_{t}$, whereas in watermarked generation, this independence is deliberately violated. Therefore, watermark detection can be formulated as a hypothesis testing problem on the dependence structure between the observed tokens and the pseudo-random variables (Aaronson, 2023; Li et al., 2025b, c; Piet et al., 2025).
Despite substantial progress, existing watermark detection methods still face several challenges. First, most procedures are designed for fixed-sample testing and lack rigorous guarantees under optional stopping. This limitation is particularly problematic in real-world implementation, where LLMs generate text as a stream. In such sequential settings, repeatedly checking for evidence of the watermark using traditional p-value methods may inflate false positive rates, yet waiting for the full text to be generated introduces high computational costs and unacceptable latency for applications requiring immediate intervention. This challenge is especially critical in the emerging context of autonomous agents; as AI tools like OpenClaw automate workflows in a streamlined manner, there is a pressing need to distinguish between human and AI-generated actions in an online fashion. [^1] As an example of increasingly autonomous agent behavior in online environments, see the discussion at https://www.reddit.com/r/singularity/comments/1r3fy5s/ai_agent_melts_down_after_github_rejection_calls/. Second, statistical power of p-value methods may deteriorate when the NTP distribution becomes highly concentrated (i.e., nearly degenerate), a phenomenon that may occur in long text generation. Third, theoretical characterizations of testing power remain limited in general watermark detection frameworks, and existing methods often lack robustness against adversaries who might attempt to corrupt only later portions of a text.
@@
To address these issues, we develop a unified framework for watermark detection based on e-values and e-processes, which are recently developed tools for online testing with useful features (Grünwald et al., 2024; Vovk and Wang, 2021; Wang and Ramdas, 2022; Ramdas et al., 2023); see Ramdas and Wang (2025) for background and a comprehensive treatment of the topic. Our framework reformulates watermark detection as sequential hypothesis testing of independence and constructs e-processes tailored to the watermark-induced structure, with a special focus on the popular Gumbel-max watermark (Aaronson, 2023). The proposed approach enjoys the following key advantages:
@@
    Under mild assumptions, the proposed methodology gives rise to the only class of admissible and unbiased e-processes, hence also the only class of admissible and unbiased sequential tests. The connection between admissible e-processes and admissible sequential tests was established by Ramdas et al. (2020).
@@
    By using adaptive weights and the online Grenander algorithm (Hore et al., 2026) to construct calibrators, the proposed e-processes are flexible and have good performance. In some situations, they can even have better performance than the existing methods that do not have sequential validity.
  4. 4.
@@
Together, these contributions provide a principled and theoretically grounded approach to watermark detection. Our framework unifies existing ideas under the lens of e-processes, clarifies the statistical structure of watermark detection, and offers both finite-sample guarantees and asymptotic efficiency. The anytime validity allows the detector to continuously monitor streaming outputs and stop as soon as sufficient evidence accumulates, thereby minimizing latency and computational cost without compromising statistical rigor.
Although our study is motivated by the LLM watermark testing problem, the proposed methods and results are generally applicable to any online testing problems where independent pivotal statistics are sequentially available. This will be clear from our assumptions, most of which only concern the distributions of the test statistics under the null and the alternative hypothesis.
This article is organized as follows. Section 2 introduces preliminary notions, including the problem setting and an overview of e-processes. In Section 3, we rigorously formulate the watermark detection problem and derive the information bound measuring the difference between the null and alternative distributions, which applies to any watermark scheme. Building upon this foundation, Section 4 establishes a unified framework for watermark detection via e-processes with detailed construction strategies. In Section 5, we provide theoretical characterizations of the proposed procedures, focusing on their power properties. Empirical evaluations on both simulated data and open-source large language models are presented in Section 6 to demonstrate the practical advantages of our framework. Finally, Section 7 concludes the paper and discusses potential avenues for future research.
@@ 3.1 Watermarks and the detection problem
and $\mathbb{Q}$ is an objective data generating probability (Li et al., 2025b). The condition in (1) can be rewritten as $S(\mathbf{P},\zeta)\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{d}}}\mathrm{Cat}(\mathbf{P})$ for each $\mathbf{P}\in\Delta_{K}$, where $\mathrm{Cat}(\mathbf{P})$ is the categorical distribution with probability vector $\mathbf{P}$. There is no other requirement on $S$.
@@
For a given $\mathbf{P}$, we use $\mathbb{U}^{\mathbf{P}}$ for the distribution of $(W,\zeta)$ under $H_{0}$ and $\mathbb{Q}^{\mathbf{P}}$ for the distribution of $(W,\zeta)$ under $H_{1}$, to emphasize that both $\mathbb{U}^{\mathbf{P}}$ and $\mathbb{Q}^{\mathbf{P}}$ depends on the usually unobservable $\mathbf{P}$. This makes the testing problem one with a composite null and a composite alternative. Moreover, $\mathbb{Q}^{\mathbf{P}}$ depends on the watermarking scheme $S$ through the condition (1) on $\mathbb{Q}^{\mathbf{P}}$.
@@ 3.3 Static information bounds
and it is $\infty$ otherwise. By Ramdas and Wang (2025, Theorem 3.20),
@@
it is necessary and sufficient that $E$ is given by $E=\mathrm{d}\mathbb{Q}^{\mathbf{P}}/\mathrm{d}\mathbb{U}^{\mathbf{P}}$ (Ramdas and Wang, 2025, Theorem 3.20), which depends on the NTP vector $\mathbf{P}$. Moreover, this is usually not a feasible choice of an e-variable, because the constraint in (2),
@@ 3.4 The Gumbel-max watermark
We next describe the Gumbel-max watermark, which is proposed by Aaronson (2023). As one of the most influential watermark schemes, it has been implemented internally at OpenAI (OpenAI, 2024) and has become a baseline in many studies (Li et al., 2025b, c). This watermark utilizes the Gumbel-max trick (Gumbel, 1954; Jang et al., 2017) to ensure the unbiasedness characterized in (1). The Gumbel-max watermark operates with the pseudo-random vector $\zeta=(U_{w})_{w\in\mathcal{W}}$ consisting of iid standard uniform variables, and the decoder is given by
@@
As suggested by Aaronson (2023), the pivotal statistic for one token is
@@
Putting the subscript $t$ back to the original detection problem for $T<\infty$ tokens, we denote by $Y_{t}=U_{t,W_{t}}$ the pivotal statistic for the $t$-th token $W_{t}$. Aaronson (2023) proposed to declare the presence of the watermark if the sum-based statistic $H_{\mathrm{ars}}=\sum_{t=1}^{T}h_{\mathrm{ars}}(Y_{t})$, with the score function $h_{\mathrm{ars}}(y)=-\log(1-y)$, is above a certain threshold. Other score functions, such as the log function $h_{\mathrm{log}}(y)=\log y$ (Fernandez et al., 2023; Kuditipudi et al., 2024), and the optimized score function $h_{\mathrm{gum}}^{*}$ analyzed in Li et al. (2025b), can also be adopted in the sum-based approach.
@@
for $x,y\in A$. We refer to Marshall et al. (2011) for a treatment of majorization and Schur-convexity.
@@
By Proposition C.1 of Marshall et al. (2011), the sum of univariate convex functions is Schur-convex. Therefore, for statement (i), it suffices to prove that $p\mapsto py^{1/p}$ is convex on $[0,1]$, which follows by straightforward calculus. For statement (ii), it suffices to note that $\mathbf{P}\preccurlyeq\mathbf{P}^{*}$ for all $\mathbf{P}\in\Delta^{\delta}_{K}$, which can be checked by definition. Statement (iii) follows directly from $F^{\mathbf{P}}(y)\leq F^{\mathbf{P}^{*}}(y)$ for any $y\in[0,1]$ and $\mathbf{P}\in\Delta_{K}^{\delta}$, as shown in (ii). ∎
Using the Gumbel-max watermark, the maximum e-power for an e-variable is the KL divergence between the two distributions (Theorem 3.20 of Ramdas and Wang, 2025), that is,
@@ 4 E-processes for watermarks
Throughout this section, we let $Y_{t},~t\in\mathbb{T}$ be the pivotal statistics for a watermark scheme. Recall that $\mathcal{F}=(\mathcal{F}_{t})_{t\in\mathbb{T}_{0}}$ is the information filtration, to which $(Y_{t})_{t\in\mathbb{T}}$ is adapted, with the interpretation that $\mathcal{F}_{t-1}$ contains all the statistical information available before generating the token $w_{t}$. The sequence $(Y_{t})_{t\in\mathbb{T}}$ satisfies two basic conditions.
@@
    under the null hypothesis, $Y_{t}$ is uniformly distributed on $[0,1]$ and independent of $\mathcal{F}_{t-1}$;
@@
    under the alternative hypothesis, $Y_{t}$ is super-uniform conditional on $\mathcal{F}_{t-1}$.
@@ 4.1 Online test with e-processes
We first introduce (p-to-e) calibrators (Vovk and Wang, 2021), which convert p-values into e-values. A calibrator is a decreasing function $f:[0,1]\to\mathbb{R}_{+}$ satisfying $\int_{0}^{1}f(p)\mathrm{d}p=1$. Although in some literature only $\int_{0}^{1}f(p)\mathrm{d}p\leq 1$ is required, it is without loss of generality to consider only those with $\int_{0}^{1}f(p)\mathrm{d}p=1$ because a large e-value is better for the test.
@@
where each $E_{t}=f_{t}(1-Y_{t})$ for some $f_{t}\in\Phi_{t}$.
Adapting the standard terminology in hypothesis testing, we say that an e-process $M$ is unbiased (against the alternative hypothesis) if for all probability measures $\mathbb{Q}$ in the alternative hypothesis, $\mathbb{E}^{\mathbb{Q}}[M_{t}]\geq 1$ for all $t\in\mathbb{T}$. In other words, its mean is at most $1$ under the null hypothesis, and at least $1$ under the alternative hypothesis.
The next result, which is the most important technical tool for constructing our tests, follows from standard arguments in the e-value literature.
@@
Under Assumption 1, the process $M=(M_{t})_{t\in\mathbb{T}_{0}}$ in (9) with
@@
is an unbiased e-process (also a martingale), where $f_{t}\in\Phi_{t}$ for each $t$. In particular, under the null hypothesis,
@@
$\mathbb{E}^{\mathbb{P}}[E_{t}|\mathcal{F}_{t-1}]=\mathbb{E}^{\mathbb{P}}[f_{t}(1-Y_{t})|\mathcal{F}_{t-1}]=\int_{0}^{1}f_{t}(p)\mathrm{d}p=1,$
@@
where we have used that $Y_{t}$ is independent of $\mathcal{F}_{t-1}$ and that $f_{t}$ is determined by $Y_{1},\dots,Y_{t-1}$ and it integrates to $1$. Moreover, it is clear that $E_{t}$ is nonnegative. Therefore,
$\mathbb{E}^{\mathbb{P}}[M_{t}|\mathcal{F}_{t-1}]=M_{t-1},~~~t\in\mathbb{T}.$
@@
Thus, $M$ is a nonnegative martingale starting at $1$ and hence an e-process. The probability guarantee follows from Ville’s inequality or the optional stopping theorem.
For the last statement, it suffices to note that, for each $\mathbb{Q}$ in the alternative hypothesis, using the superuniformity of $Y_{t}$ and the fact that $f_{t}$ is decreasing, we have $\mathbb{E}^{\mathbb{Q}}[f_{t}(1-Y_{t})|\mathcal{F}_{t-1}]\geq 1$ for all $t\in\mathbb{T}$. Therefore, $\mathbb{E}^{\mathbb{Q}}[M_{t}]\geq 1$. ∎
The next result gives a converse statement to Theorem 1, showing that (9) is the only possible form for e-processes when $\mathcal{F}$ is the natural filtration of $(Y_{t})_{t\in\mathbb{T}}$. An e-process $M$ is admissible if there is no other e-process $M^{\prime}$ such that $M^{\prime}\geq M$ with $\mathbb{P}(M^{\prime}_{t}>M_{t})>0$ for some $t\in\mathbb{T}$.
Theorem 2.
Suppose that $\mathcal{F}_{t}=\sigma(Y_{1},\dots,Y_{t})$ for $t\in\mathbb{T}$ and Assumption 1 holds. Then $M$ is an admissible and unbiased e-process if and only if
$M_{t}=\prod_{s=1}^{t}f_{s}(1-Y_{s}),~~~t\in\mathbb{T}_{0},~\mbox{~$\mathbb{P}$-almost surely}.$ (10)
for some $f_{t}\in\Phi_{t}$ for all $t\in\mathbb{T}$.
Proof.
The “if” statement follows from Theorem 1, and we show the “only if” statement. By Ramdas et al. (2020, Theorem 18), an admissible e-process $M$ is a $\mathbb{P}$-nonnegative martingale with $M_{0}=1$. Let $E_{t}=M_{t}/M_{t-1}$ with the convention $0/0=1$. Note that $\mathcal{F}_{t}=\sigma(Y_{1},\dots,Y_{t})$, and hence $E_{t}$ can be written as a function $e_{t}$ of $(Y_{1},\dots,Y_{t})$. Since $M^{\prime}$ is a nonnegative martingale and $Y_{1},\dots,Y_{t}$ are independent, we have
$\mathbb{E}^{\mathbb{P}}[e_{t}(Y_{1},\dots,Y_{t})|\mathcal{F}_{t-1}]=\int_{0}^{1}e_{t}(Y_{1},\dots,Y_{t-1},y)\mathrm{d}y=1.$
Let $f_{t}:y\mapsto e_{t}(Y_{1},\dots,Y_{t-1},1-y)$ for $t\in\mathbb{T}$, which is Borel measurable in $y$. We next show, for each $t$,
$\mathbb{P}\left(\{\mbox{$f_{t}$ is decreasing almost everywhere on $(0,1)$}\}\right)=1.$
The intuition is that if this does not hold true, then a larger $Y_{t}$ may not lead to a larger $E_{t}$, which yields $\mathbb{E}^{\mathbb{Q}}[M_{t}]<1$ for some $\mathbb{Q}$ in the alternative hypothesis.
Suppose for contradiction that there exists $t\in\mathbb{T}$ such that the set
$A:=\{\mbox{$f_{t}$ is not decreasing almost everywhere on $(0,1)$}\}\in\mathcal{F}_{t-1}$
has positive probability. For every value of $\mathbf{y}\in[0,1]^{t-1}$ taken by $\mathbf{Y}:=(Y_{1},\dots,Y_{t-1})$ on the event $A$, there exists $x^{\mathbf{y}}\in(0,1)$ such that the set
$B^{\mathbf{y}}:=\{y\in(x_{0},1):e_{t}(\mathbf{y},y)>e_{t}(\mathbf{y},x^{\mathbf{y}})\}$
has positive Lebesgue measure.
We will now construct a probability measure $\mathbb{Q}$ in the alternative hypothesis through the distribution of $Y_{t}$ under $\mathbb{Q}$. Comparing $\mathbb{P}$ and $\mathbb{Q}$, we will only modify the distribution of $Y_{t}$ on the event $A$, and everything stays unchanged outside $A$. Let $U$ be a uniform random variable on $[0,1]$ independent of everything else. Define the function $h^{\mathbf{y}}(u)=x_{0}\mathds{1}_{\{u\in B^{\mathbf{y}}\}}+u\mathds{1}_{\{u\not\in B^{\mathbf{y}}\}}$ for $y\in(0,1)$. It is clear that $h^{\mathbf{y}}(u)\leq u$ for $u\in(0,1)$ and hence $1-h^{\mathbf{y}}(U)$ is superuniform.
Under $\mathbb{Q}$, we let $Y_{t}$ be distributed as $1-h^{\mathbf{Y}}(U)$ on event $A$, which is superuniform; $(Y_{s})_{s\in\mathbb{T}}$ remains an independent sequence with $Y_{s}$ uniformly distributed for $s\neq t$. The measurability of $Y_{t}$ (to be well-defined as a random variable) is guaranteed by the assumption of independence. Because $B^{\mathbf{y}}$ has positive Lebesgue measure, $e_{t}(\mathbf{y},y)>e_{t}(\mathbf{y},x^{\mathbf{y}})$ for $y\in B^{\mathbf{y}}$ and $\int_{0}^{1}e_{t}(\mathbf{y},u)\mathrm{d}u=1$, we have
$\int_{0}^{1}e_{t}(\mathbf{y},h^{\mathbf{y}}(u))\mathrm{d}u<\int_{0}^{1}e_{t}(\mathbf{y},u)\mathrm{d}u=1.$
Therefore,
$\mathbb{E}^{\mathbb{Q}}[f_{t}(1-Y_{t})|\mathcal{F}_{t-1}]$ $=\mathbb{E}^{\mathbb{Q}}[f_{t}(h^{\mathbf{Y}}(U))|\mathcal{F}_{t-1}]\mathds{1}_{A}+\mathbb{E}^{\mathbb{Q}}[f_{t}(U)|\mathcal{F}_{t-1}]\mathds{1}_{A^{c}}$
$=\left(\int_{0}^{1}e_{t}(\mathbf{Y},h^{\mathbf{Y}}(u))\mathrm{d}u\right)\mathds{1}_{A}+\mathds{1}_{A^{c}}.$
Since $Z_{t}:=\int_{0}^{1}e_{t}(\mathbf{Y},h^{\mathbf{Y}}(u))\mathrm{d}u<1$ on $A$ and $A$ has positive probability under both $\mathbb{P}$ and $\mathbb{Q}$, we obtain
$\mathbb{E}^{\mathbb{Q}}[M_{t}]$ $=\mathbb{E}^{\mathbb{Q}}\left[M_{t-1}\mathbb{E}^{\mathbb{Q}}[f_{t}(h^{\mathbf{Y}}(U))|\mathcal{F}_{t-1}]\right]$
$=\mathbb{E}^{\mathbb{Q}}\left[M_{t-1}Z_{t}\mathds{1}_{A}+M_{t-1}\mathds{1}_{A^{c}}\right]<\mathbb{E}^{\mathbb{Q}}[M_{t-1}]=1.$
Therefore, $\mathbb{E}[M_{t}]<1$, a contradiction. Therefore, $A$ must have zero probability, and (10) holds with $f_{t}\in\Phi_{t}$ for each $t$. ∎
As a further step, Ramdas et al. (2020, Theorem 18) also implies that all admissible sequential test $(\psi_{t})_{t\in\mathbb{T}}$ at level $\alpha$ for the null hypothesis must have the form
$\psi_{t}=\max_{s\leq t}\mathds{1}_{\{M_{s}\geq 1/\alpha\}},~~~t\in\mathbb{T}$
for some e-process $M$, that is, rejecting the null hypothesis when $M$ reaches $1/\alpha$. Therefore, from Theorems 1 and 2, we conclude that the method we proposed in Algorithm 1 is essentially the only way to build sequential tests for this problem.
@@
In the next subsections, we present a few specific ways of constructing e-processes in the proposed detection procedure.
Remark 1.
It should be clear from Assumption 1 and Theorem 1 that the proposed methodology is applicable to any online testing problem where pivotal statistics are sequentially available, such that they have known continuous distributions under the null hypothesis. The uniqueness result in Theorem 2 further requires that their distributions become stochastically larger under the alternative hypothesis.